Laurel Enterprises expects earnings next year of ​$3.94 per share and has a 30% retention​ rate, which it plans to keep constant. Its equity cost of capital is 9%​, which is also its expected return on new investment. Its earnings are expected to grow forever at a rate of 2.7 % per year.

If its next dividend is due in one​ year, what do you estimate the​ firm's current stock price to​ be?
Answer #1

Year 1 dividend = EPS * Payout ratio

Year 1 dividend = 3.94 * (1 - 30%)

Year 1 dividend = 3.94 * (1 - 0.3)

Year 1 dividend = 2.758

Current stock price = D1 / required rate - growth rate

Current stock price = 2.758 / 0.09 - 0.027

Current stock price = 2.758 / 0.063

Current stock price = $43.78